OpenStudy (anonymous):

write an equation of the line with the slope m=-3 that passes through the point (2,3)

OpenStudy (anonymous):

y = mx + b plug in x, y, and m and solve for b 3 = -3(2) + b 3=-6 + b 9=b

OpenStudy (anonymous):

plug m and b into slope-intercept form y=-3x + 9

OpenStudy (anonymous):

check: 3=-3(2)+9 3=-6+9 3=3

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

Or you can just plug it into point slope.. \[y-3 = -3(x-2)\]

OpenStudy (anonymous):

welcome...great to know both ways!